# HumanVBench: Exploring Human-Centric Video Understanding Capabilities of MLLMs with Synthetic Benchmark Data

### v1

@@ Abstract
In the domain of Multimodal Large Language Models (MLLMs), achieving human-centric video understanding remains a formidable challenge. Existing benchmarks primarily emphasize object and action recognition, often neglecting the intricate nuances of human emotions, behaviors, and speech-visual alignment within video content. We present HumanVBench, an innovative benchmark meticulously crafted to bridge these gaps in the evaluation of video MLLMs. HumanVBench comprises 17 carefully designed tasks that explore two primary dimensions: inner emotion and outer manifestations, spanning static and dynamic, basic and complex, as well as single-modal and cross-modal aspects. With two advanced automated pipelines for video annotation and distractor-included QA generation, HumanVBench utilizes diverse state-of-the-art (SOTA) techniques to streamline benchmark data synthesis and quality assessment, minimizing human annotation dependency tailored to human-centric multimodal attributes. A comprehensive evaluation across 16 SOTA video MLLMs reveals notable limitations in current performance, especially in cross-modal and temporal alignment, underscoring the necessity for further refinement toward achieving more human-like understanding. HumanVBench is open-sourced to facilitate future advancements and real-world applications in video MLLMs.
Figure 1: Overview of HumanVBench, which encompasses 17 fine-grained tasks for extensive human-centric evaluations (middle blue box). Each task is denoted by its acronym and the number of included QA instances. The right orange box illustrates some examples of these QAs. HumanVBench is constructed using the novel automated Video Annotation Pipeline (upper left, purple box), followed by the Distractor-Included QA Synthesis Pipeline (lower left, green box). These pipelines are reusable and backed by more than twenty data processing operators with advanced algorithm implementation and cutting-edge auxiliary models.
@@ 1 Introduction
To bridge these gaps, we introduce HumanVBench, a pioneering benchmark specifically tailored for video MLLMs focusing on human-centric analysis. HumanVBench includes 17 meticulously designed QA tasks, categorized into two core aspects: inner emotion and outer manifestation, as illustrated in Figure 1. The inner emotion dimension evaluates the model’s capacity to perceive emotional cues from videos, while the outer manifestation encompasses person recognition, human behavior analysis, and cross-modal speech-visual alignment. Each category is further subdivided into many fine-grained tasks, offering a comprehensive evaluation across static and dynamic, basic and complex, single-modal and cross-modal aspects. For instance, it enables us to examine MLLMs’ understanding of human actions, emotions, and the associated causal relationships, as well as considering the alignment degree between appearance, mouth movement, audio, and speech content in videos.
@@
Through HumanVBench, we comprehensively evaluate 15 SOTA video MLLMs, including open-source models like LLaVA-OneVision [22] and commercial ones like GPT-4o [35]. Our evaluations reveal several interesting insights and significant gaps between current model capabilities and human-like understanding, particularly in tasks such as cross-modal alignment and emotion perception. While proprietary models demonstrate closer human-like accuracy, open-source models frequently misclassify emotions due to temporal noise, underscoring the need for further architectural improvements and refined datasets.
@@
  • We open source our benchmark, including data, evaluation and synthesis codes at https://github.com/modelscope/data-juicer/tree/HumanVBench, to shield light on further evolution of future human-centric video analysis systems.
@@ 2 Related Works
Multimodal Large Language Models.
The remarkable progress in Large Language Models (LLMs) has sparked extensive research into merging language comprehension with visual and auditory information, thereby expediting the advancement of multimodal models. Within this domain, image MLLMs amalgamate visual and linguistic data to enhance image interpretation and cross-modal reasoning [9, 26, 17, 31]. Video MLLMs [23, 33, 28, 42, 51, 10] extend these capabilities by incorporating temporal sequences for dynamic video analysis. Furthermore, generalist MLLMs [14, 13, 54] are designed to process a plethora of inputs, including images, video, and audio, thus improving adaptive task performance across various modalities. Despite these advancements, rigorous evaluation of video MLLMs on human-centric video understanding tasks remains an unaddressed challenge, which this work aims to address.
Video Benchmarks for MLLMs.
There exists a substantial array of benchmarks for assessing the performance of MLLMs, with a particular emphasis on video-based evaluations. Presently, general-purpose video benchmarks can be broadly categorized based on their evaluative methodologies: multiple-choice queries with definitive answers [34, 36, 24, 12], Open-Ended Video Question Answering necessitating supplementary LLMs for evaluation [46, 33, 11], video-captioning benchmarks [46, 8], and video-generation assessments such as VBench [16]. Additionally, there are human-centric benchmarks in MLLMs. ActivityNet-QA [46], rooted in the widely-used ActivityNet dataset [50], emphasizes object and human activity recognition. HumanBench [40] amalgamates multiple public datasets to create benchmarks for tasks like pose estimation, pedestrian attribute recognition, and crowd assessment. HERM [25] introduces a human-centric image benchmark focusing on people, poses, actions, and interactions, using GPT-4-generated question-answer pairs.
Our Position.
In comparison with existing benchmarks, HumanVBench distinguishes itself through both innovative task design and unique construction methodologies. Firstly, it introduces novel human-related video tasks. Unlike HERM, HumanVBench integrates complex scenarios encompassing temporal dynamics for enhanced human video interpretation; compared to ActivityNet-QA and HumanBench, HumanVBench delves into nuanced emotional comprehension and evaluates cross-modal alignment between character visual data and speech modalities. Moreover, HumanVBench boasts a pioneering construction approach, deriving from raw, uncurated video content and employing a suite of over 20 advanced processing operators for meticulous video character annotation and automatized multiple-choice question formulation. Human intervention is minimized to verifying the quality of well-structured QAs.
@@ 3.1 Task Design and Definition
Human observers, by nature, focus intently on individuals within videos, analyzing their appearance, emotional states, and behaviors. This inherent focus forms the basis of our systematic design of 17 fine-grained human-related tasks intended to assess the ability of MLLMs to emulate human-like perception and comprehension in video analysis (Figure 1). Detailed definitions and examples for each task are available in the Appendix. These tasks are categorized based on the observability of analyzed content into two categories: Inner Emotion and Outer Manifestation.
@@ 3.1.1 Inner Emotion
In realistic video scenarios, inner emotions are often less observable due to the possible lack of focal points on a person’s face, as cameras might capture a wider array of scenes. Thus, emotion perception becomes an abstract skill requiring careful attention to facial expressions, body language, and verbal cues. With this goal, tasks designed under this category assess the capability of video-MLLMs to capture and interpret emotional subtleties akin to human perceptual abilities.
@@
  • Emotion Recognition in Conversation expands on previous tasks by integrating speech content from the MELD dataset [38] to judge multimodal emotional comprehension when text is available.
@@ 3.1.2 Outer Manifestation
  • Human Counting (HC) investigates the model’s proficiency in detecting, remembering, and counting distinct individuals in a video scene.
  • Appearance Time Detection (ATD) requires accurate identification of a specified individual’s presence time in the video.
Besides, we incorporate the Human Behavior Analysis category to scrutinize the model’s abilities for understanding and analyzing individual behaviors. It includes four tasks:
@@
  • Audio-Visual Speaker Matching (AVSM) involves correlating audio features to identify individuals, and analyzing potential matches in domains such as appearance, gender, and age.
@@
These aforementioned tasks together enable a comprehensive assessment of MLLMs’ diverse perceptual abilities across emotion, identity, behavior and speech-visual alignment, driving advancements toward human-like video comprehension.
@@ 3.2 Human-Centric Video Annotation Pipeline
Creating task-specific questions for the aforementioned tasks hinges on extensive human-centric annotations within videos. Our video annotation pipeline, illustrated in Figure 2, emphasizes multi-modal, granular annotation pathways, supported by the Data-Juicer [4, 6] framework. While benefiting from existing operators, we’ve additionally developed innovative operators to contribute back to the open-source community. Detailed annotations and examples are available in the Appendix.
@@ 3.2.1 Collecting Videos Containing People
We sourced copyright-free videos from Pexels [37], applying the video_split_by_scene_mapper to partition videos based on scene transitions of ensuring accurate human tracking. After a series of attribute-based filtering such as duration and optical flow, a video_face_ratio_filter gauges the visibility of faces, retaining videos where a face appears in a majority of the frames for further subsequent annotation.
@@ 3.2.2 Video Mappers
First, we implement a specially designed operator named video_human_tracks_extraction_mapper to track frame-level face bounding box, which is essential for subsequent person localization and highlighting. A track is constructed by iteratively detecting the bounding boxes across frames based on thresholds for sequential frame overlaps, with the aim to ensure accurate person localization. This process also allows for an approximate count of how many people appear within the continuous shot.
Next, the human_demographics_mapper is leveraged to assign demographic labels such as age, gender, and race. Bounding box data at different positions such as full-body or face-only enable flexible cropping to generate individual-focused videos, paying attention to appearance and posture (via the video_human_description_mapper) or facial expressions and transformations (via the video_facial_description_mapper). Several distinct benefits were derived from these processes. On the one hand, it prevents the description model from failing to recognize videos with bounding boxes or being distracted by other people in the frame. On the other hand, it helps preserve facial details that could be lost through resizing.
Further, a series of audio operations are employed to enrich information about people in the video. The audio annotation process begins with a audio_tagging_mapper to classify the sound type. For videos whose sound type is determined to be “speech”, specific operators are activated. These include the active_speaker_detection_mapper for identifying who is the real speaker in the video, the asr_mapper for transcribing spoken content, the speech_emotion_recognition_mapper for detecting emotional cues in speech, and the voice_demographics_mapper for profiling voice characteristics, such as gender and age.
Finally, annotations will be extended to encompass event atmosphere and broader video narratives, captured via video_description_mapper. As a result, these detailed annotations enable versatile and automated question synthesis. For instance, in the case of Audio-Visual Speaker Matching task, videos are filtered using speech features (Label-10 and Label-3 of Figure 2) for precise alignment against tracked individuals (Label-1 becomes the basis for aligning detected features, thus determining ground truths). The complete construction details of all tasks is presented in the Appendix.
@@ 3.3 Distractor-Included QA Generation Pipeline
For tasks with exact answers (e.g. restricted category, number or letter), their questions, correct answers, and distractors are constructed using specially designed templates and annotations from section 3.2. For open-ended questions, we designed the following pipeline to construct questions, answers, and distractors. These tasks include Human Emotion Recognition, Emotion Temporal Analysis, Emotion Intensity Comparison, Human-to-Text, Behavior Temporal Analysis, and Behavior Causality Analysis. The Distractor-Included QA Synthesis Pipeline is shown in Figure 3.
@@ 3.3.1 Selecting Videos for Generating Questions
For a specific task, it is first necessary to filter annotated videos to select those suitable for question generation. This can be done based on video length and the number of people in the video. For example, the Behavior Temporal Analysis task may require longer videos than other tasks, and Human-to-Text task requires videos with at least two people. After filtering by these criteria, a subset of videos suitable for question generation is obtained.
Table 1: The performance of SOTA video MLLMs on HumanVBench, grouped into 9 visual-only MLLMs, 5 audio-visual MLLMs and 2 proprietary ones. “Random” represents the results of random guessing, while “Human” represents human-level performance. The task acronym can be refered in Figure 1. For each task, the best result across all tested models is bolded, and the best result among open-source models is underlined. “-” means the model does not have all the modal processing capabilities of the input required for this task. More evaluation implementation details are presented in Appendix.
@@
Models Frames ER ETA AR EIC ERC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 25.0 25.0 25.0 25.0 14.3 22.9 27.9 25.0 23.1 25.0 25.3 25.0 25.0 25.0 20.0 23.8 42.8 23.6 33.3 25.2 31.2
Chat-UniVi 1 f/s 40.6 19.9 10.9 16.3 26.1 22.8 29.1 52.6 17.8 19.8 29.8 28.4 44.1 7.4 6.7 21.7 - - - - -
CogVLM2-Video 1 f/s 49.1 33.1 36.0 11.1 32.4 32.3 42.5 68.5 31.1 40.0 45.5 52.9 63.8 40.0 14.1 42.7 - - - - -
VILA 4 f 47.2 34.4 28.7 14.7 34.1 31.8 50.4 62.4 40.4 20.2 43.4 62.4 66.2 33.1 51.4 53.3 - - - - -
Video-LLaVA 8 f 42.5 25.7 26.2 24.7 33.8 30.6 27.9 64.1 28.3 31.7 38.0 44.1 55.3 34.3 40.7 43.6 - - - - -
LLaVAOneVision 8 f 57.5 33.1 62.8 15.7 42.9 42.4 67.0 73.9 49.1 10.0 50.0 72.4 74.5 37.1 47.5 57.9 - - - - -
PLLaVA 16 f 41.9 24.3 36.6 23.2 45.7 34.3 33.5 67.4 29.9 23.3 38.5 42.4 53.2 40.0 14.7 37.6 - - - - -
ShareGPT4Video 16 f 55.1 30.9 40.9 10.1 39.0 35.2 33.0 56.5 24.5 43.3 39.3 52.9 62.4 41.4 16.9 43.4 - - - - -
Otter-V 16 f 19.8 25.7 26.2 21.2 12.9 21.2 30.2 23.9 20.8 22.5 24.4 24.7 20.6 25.7 23.1 23.5 - - - - -
VideoChat2-IT 16 f 27.5 30.9 39.6 25.3 32.9 31.2 20.1 51.1 11.2 26.7 27.3 38.8 41.1 34.3 24.9 34.8 - - - - -
Video-LLaMA 8 f 34.7 25.9 11.5 19.7 25.0 23.4 28.4 28.5 23.9 26.5 26.8 35.5 35.9 22.6 15.4 27.4 40.1 26.6 33.1 26.2 31.5
Video-LLaMA2.1 16 f 49.1 28.7 47.0 21.7 36.7 36.6 34.6 58.7 41.5 16.7 37.9 53.5 65.2 37.1 33.3 47.3 44.0 27.1 27.7 23.7 30.6
ImageBind-LLM 15 f 25.1 21.0 25.0 26.4 20.8 23.7 23.5 25.9 23.8 19.5 23.2 25.3 30.4 31.4 23.3 27.6 45.0 25.1 28.9 22.9 30.5
ChatBridge 4 f 38.3 15.6 30.2 14.4 27.4 25.2 31.4 58.9 23.7 8.7 30.7 38.0 51.3 33.1 16.9 34.8 43.7 25.3 30.4 34.1 33.4
OneLLM 15 f 41.7 29.4 34.1 21.1 29.8 31.2 29.0 50.4 20.6 27.8 32.0 32.9 57.7 25.7 22.4 34.7 43.4 26.4 29.5 23.2 30.6
GPT-4o 60.5 38.2 12.2 64.1 48.1 44.6 46.9 84.8 47.2 25.8 51.2 83.5 78.0 47.1 61.6 67.6 - - - - -
Gemini-1.5-Pro 67.1 54.4 53.0 67.2 55.2 59.4 87.1 82.6 52.8 71.7 73.6 85.9 78.0 64.3 75.1 75.8 90.1 76.8 66.4 88.5 80.5
Human 91.6 86.0 87.8 80.8 56.0 80.4 98.9 85.8 92.5 78.3 88.9 93.0 86.5 88.6 88.1 89.1 96.0 96.1 87.0 88.0 91.8
Figure 4: Two examples of 8 frames sampled at equal intervals from a video of a speaker.
Table 2: Video-MLLMs accuracy on Human Emotion Recognition (HER) and Speaker Emotion Recognition (SER) tasks. Current video video MLLMs struggle with speaker-centric tasks within HumanVBench.
Visual-only Video-MLLMs Audio-video MLLMs
Models Chat-UniVi Cog-VLM2-Video VILA Video-LLaVA LLAVA-One-Vision PLLaVA Share-GPT4-Video GPT-4o Avg. Dec. Chat-Bridge One-LLM Video-LLAMA Video-LLAMA-2.1 Gemini-1.5-Pro Avg. Dec.
Emo. Acc. 40.6 49.1 47.2 42.5 57.5 41.9 55.1 60.5 49.3 38.3 41.7 34.7 49.1 67.1 46.2
Speaker Emo. Acc. 36.3$\downarrow$-4.3 40.2$\downarrow$-8.9 41.3$\downarrow$-5.9 39.3$\downarrow$-3.2 51.8$\downarrow$-5.7 35.7$\downarrow$-6.2 47.3$\downarrow$-7.8 53.6$\downarrow$-6.9 $\downarrow$-6.1 33.4$\downarrow$-4.9 39.8$\downarrow$-1.9 33.0$\downarrow$-1.7 42.0$\downarrow$-7.1 60.7$\downarrow$-6.4 $\downarrow$-4.4
Figure 5: Distribution of emotion type recognition capability. (a) Accuracy distribution for Emotion Recognition in Conversation tasks by emotion type. (b) Accuracy distribution for human emotion recognition tasks by emotive polarity.
3.3.2 Synthesizing Questions and Preliminary Answers

### v2

@@ Abstract
In the domain of Multimodal Large Language Models (MLLMs), achieving human-centric video understanding remains a formidable challenge. Existing benchmarks primarily emphasize object and action recognition, often neglecting the intricate nuances of human emotions, behaviors, and speech-visual alignment within video content. We present HumanVBench, an innovative benchmark meticulously crafted to bridge these gaps in the evaluation of video MLLMs. HumanVBench comprises 16 carefully designed tasks that explore two primary dimensions: inner emotion and outer manifestations, spanning static and dynamic, basic and complex, as well as single-modal and cross-modal aspects. With two advanced automated pipelines for video annotation and distractor-included QA generation, HumanVBench utilizes diverse state-of-the-art (SOTA) techniques to streamline benchmark data synthesis and quality assessment, minimizing human annotation dependency tailored to human-centric multimodal attributes. A comprehensive evaluation across 22 SOTA video MLLMs reveals notable limitations in current performance, especially in cross-modal and emotion perception, underscoring the necessity for further refinement toward achieving more human-like understanding. HumanVBench is open-sourced to facilitate future advancements and real-world applications in video MLLMs.
Figure 1: Overview of HumanVBench, which encompasses 16 fine-grained tasks for extensive human-centric evaluations (middle blue box). Each task is denoted by its acronym and the number of included QA instances. The right orange box illustrates some examples of these QAs. HumanVBench is constructed using the novel automated Video Annotation Pipeline (upper left, purple box), followed by the Distractor-Included QA Synthesis Pipeline (lower left, green box). These pipelines are reusable and backed by more than twenty data processing operators with advanced algorithm implementation and cutting-edge auxiliary models.
@@ 1 Introduction
To bridge these gaps, we introduce HumanVBench, a pioneering benchmark specifically tailored for video MLLMs focusing on human-centric analysis. HumanVBench includes 16 meticulously designed QA tasks, categorized into two core aspects: inner emotion and outer manifestation, as illustrated in Figure 1. The inner emotion dimension evaluates the model’s capacity to perceive emotional cues from videos, while the outer manifestation encompasses person recognition, human behavior analysis, and cross-modal speech-visual alignment. Each category is further subdivided into many fine-grained tasks, offering a comprehensive evaluation across static and dynamic, basic and complex, single-modal and cross-modal aspects. For instance, it enables us to examine MLLMs’ understanding of human actions, emotions, and the associated causal relationships, as well as consider the alignment degree between appearance, mouth movement, audio, and speech content in videos.
@@
Through HumanVBench, we comprehensively evaluate 22 SOTA video MLLMs, including open-source models like VideoLLaMA3 [57] and commercial ones like GPT-4o [41]. Our evaluations reveal several interesting insights and significant gaps between current model capabilities and human-like understanding, particularly in tasks such as cross-modal alignment and emotion perception. While proprietary models demonstrate closer human-like accuracy, open-source models frequently misclassify emotions due to temporal noise, underscoring the need for further architectural improvements and refined datasets.
@@
  • We release our benchmark, including data, evaluation, and synthesis codes at https://github.com/modelscope/data-juicer/tree/HumanVBench , to foster further evolution of future human-centric video analysis systems.
@@ 2 Related Works
Multimodal Large Language Models. The remarkable progress in Large Language Models (LLMs) has sparked extensive research into merging language comprehension with visual and auditory information, thereby expediting the advancement of multimodal models. Within this domain, image MLLMs amalgamate visual and linguistic data to enhance image interpretation and cross-modal reasoning [13, 32, 23, 37]. Video MLLMs [29, 39, 34, 48, 58, 16] extend these capabilities by incorporating temporal sequences for dynamic video analysis. Furthermore, generalist MLLMs [20, 19, 61] can process a plethora of inputs, including images, video, and audio, thus improving adaptive task performance across various modalities. Despite these advancements, rigorous evaluation of video MLLMs on human-centric video understanding tasks remains an unaddressed challenge, which this work aims to address.
Video Benchmarks for MLLMs. There exists a substantial array of benchmarks for assessing the performance of MLLMs, with a particular emphasis on video-based evaluations. Presently, general-purpose video benchmarks can be broadly categorized based on their evaluative methodologies: multiple-choice queries with definitive answers [40, 42, 30, 18, 45], Open-Ended Video Question Answering necessitating supplementary LLMs for evaluation [52, 39, 17, 45], video-captioning benchmarks [52, 12, 5], and video-generation assessments such as VBench [22]. Additionally, there are human-centric benchmarks in MLLMs. ActivityNet-QA [52], rooted in the widely-used ActivityNet dataset [56], emphasizes object and human activity recognition. HumanBench [46] amalgamates multiple public datasets to create benchmarks for tasks like pose estimation, pedestrian attribute recognition, and crowd assessment. HERM [31] introduces a human-centric image benchmark focusing on people, poses, actions, and interactions, using GPT-4-generated question-answer pairs.
Our Position. Compared to existing benchmarks, HumanVBench distinguishes itself through both innovative task design and unique construction methodologies. Firstly, it introduces novel human-related video tasks. Unlike HERM, HumanVBench integrates complex scenarios encompassing temporal dynamics for enhanced human video interpretation; compared to ActivityNet-QA and HumanBench, HumanVBench delves into nuanced emotional comprehension and evaluates cross-modal alignment between visual data and speech modalities. Moreover, HumanVBench boasts a pioneering construction approach, deriving from raw, uncurated video content and employing over 20 advanced processing operators for meticulous video character annotation and automatized multiple-choice question formulation. Human intervention is minimized to verify the quality of well-structured QAs.
@@ 3.1 Task Design and Definition
Human observers naturally concentrate on individuals in videos, examining their appearance, emotions, and behaviors. This intrinsic focus underpins our design of 16 fine-grained, human-centric tasks, aimed at evaluating MLLMs’ ability to mimic human-like perception and understanding in video analysis (Figure 1). Each task is detailed with definitions and examples in the Appendix Section 10. The tasks are grouped into two categories based on content observability: Inner Emotion and Outer Manifestation.
@@ 3.1.1 Inner Emotion
In real-world videos, inner emotions are less observable, as cameras often capture wider scenes instead of facial close-ups. Thus, emotion perception becomes an abstract skill requiring careful attention to facial expressions, body language, and verbal cues. With this goal, tasks in this category evaluate video-MLLMs’ ability to detect and interpret emotional nuances, mirroring human perceptual skills.
@@ 3.1.2 Outer Manifestation
  • Human Counting (HC) evaluates the model’s ability to detect, track, and count distinct individuals in a video.
  • Appearance Time Detection (ATD) requires identifying a specified individual’s presence time in the video.
Besides, we incorporate the Human Behavior Analysis category to scrutinize the model’s ability to understand and analyze individual behaviors, with four tasks:
@@
  • Audio-Visual Speaker Matching (AVSM) correlates audio features to identify individuals and analyzes matches by appearance (gender and age).
@@
These aforementioned tasks together enable a comprehensive assessment of MLLMs’ diverse perceptual abilities across emotion, identity, behavior and speech-visual alignment, pushing toward human-like video comprehension.
@@ 3.2 Human-Centric Video Annotation Pipeline
Creating task-specific questions for the aforementioned tasks hinges on extensive human-centric annotations within videos. Our video annotation pipeline, illustrated in Figure 2, emphasizes multi-modal, granular annotation pathways, supported by the Data-Juicer [6, 9] framework. While benefiting from existing operators, we’ve also developed innovative operators to enhance the open-source community. Detailed annotations and examples are in the Appendix Section 11.
@@ 3.2.1 Collecting Videos Containing People
We sourced copyright-free videos from Pexels [43], splitting each video by scene transitions using video_split_by_scene_mapper for accurate human tracking. After a series of attribute-based filtering such as duration and optical flow, a video_face_ratio_filter gauges the visibility of faces, retaining videos where a face appears in most frames for further subsequent annotation.
@@ 3.2.2 Video Mappers
We begin by implementing a specially designed operator named video_human_tracks_extraction_mapper to track each person across consecutive frames. The operator constructs tracks by iteratively detecting face and body bounding boxes, using overlap thresholds between consecutive frames to ensure accurate and consistent person localization. Additionally, this process enables an approximate count of the number of people appearing within the continuous shot. The generated tracks serve as a foundation for subsequent tasks, including person highlighting and detailed descriptions of their appearances or actions.
Next, visual attributes and descriptions for each individual are derived from track information. Specifically, the human_demographics_mapper extracts facial images from face bounding box tracks and employs existing facial attribute detection models to infer demographic labels (e.g., age, gender, race). Additionally, bounding box data at various positions (e.g., full-body or face-only) enables flexible cropping to generate individual-focused videos. This allows the description model to concentrate on appearance and posture (via the video_human_description_mapper) or facial expressions and transformations (via the video_facial_description_mapper). These processes effectively prevent the description model from misidentifying bounding box contexts or being disrupted by other individuals in the frame, while ensuring facial details are preserved during resizing.
Further, audio operations enrich video content by analyzing people-related information. The annotation process initiates with the audio_tagging_mapper classifying sound types. For speech-detected videos, specialized operators are activated: the active_speaker_detection_mapper identifies active speakers by fusing audio and visual cues, the asr_mapper transcribes speech content, the speech_emotion_recognition_mapper extracts emotional cues, and the voice_demographics_mapper profiles voice characteristics (e.g., gender, age).
Finally, annotations are extended to capture event atmosphere and broader video narratives through the video_description_mapper. These detailed annotations facilitate versatile, automated question synthesis. For example, in the Audio-Visual Speaker Matching task construction, videos are filtered by verifying that exactly one individual’s Label-3 (visual demographics) matches the video’s Label-10 (audio demographics), with all other individuals exhibiting mismatches. The matching individual is established as the ground truth, with others marked as interference, while Label-1 tracking data assigns distinct identifiers to each person. Complete task construction details are provided in the Appendix Section 12.
@@ 3.3 Distractor-Included QA Generation Pipeline
For tasks with exact answers (e.g., restricted categories, numbers, or letters), questions, correct answers, and distractors are constructed using tailored templates and annotations from Section 3.2. For open-ended questions, we developed a pipeline to generate questions, answers, and distractors, applied to six tasks in HumanVBench: Human Emotion Recognition, Emotion Temporal Analysis, Emotion Intensity Comparison, Human-to-Text, Behavior Temporal Analysis, and Behavior Causality Analysis. The Distractor-Included QA Synthesis Pipeline is illustrated in Figure 3.
@@ 3.3.1 Selecting Videos for Generating Questions
For a specific task, annotated videos are filtered based on task-specific criteria, such as video length and the number of people, to select suitable candidates for question generation. For instance, the Behavior Temporal Analysis task requires longer videos, while Human-to-Text necessitates at least two people. This process yields a subset of videos tailored for question generation.
@@
Models Frames ER ETA AR EIC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 25 25.0 25.0 25.0 22.9 27.9 25.0 23.1 25.0 25.3 25.0 25.0 25.0 20.0 23.8 42.8 23.6 33.3 25.0 31.2
Chat-UniVi 1 f/s 29.2 19.9 10.9 16.3 18.8 29.1 37.2 17.8 19.8 26.0 22.0 19.4 14.0 6.7 15.5 42.8 20.6 26.3 18.6 27.1
CogVLM2-Video 1 f/s 25.0 33.1 36.0 11.1 25.8 42.5 63.8 31.1 40.0 44.4 42.4 47.8 34.0 14.1 34.6 59.2 38.7 30.7 17.9 36.6
VideoLLaMA3 1 f/s 39.4 34.5 48.0 43.1 41.3 90.3 78.1 44.9 71.0 71.1 72.4 53.4 56.4 69.0 62.8 64.6 63.1 34.4 17.9 45.0
VILA 6 f 29.1 34.4 28.7 14.7 26.7 50.4 54.2 40.4 20.2 41.3 53.0 54.5 33.0 51.4 48.0 47.4 23.9 34.3 18.6 31.1
Video-LLaVA 8 f 27.1 25.7 26.2 24.7 25.4 27.9 46.8 28.3 31.7 33.7 34.8 37.3 27.0 40.7 35.0 50.0 28.4 34.3 21.2 33.5
LLaVAOneVision 8 f 36.5 33.1 62.8 15.7 36.0 67.0 68.1 49.1 10.0 48.6 63.6 55.2 37.0 47.5 50.8 52.6 51.6 35.0 26.9 42.8
InternVL2 8 f 35.4 29.3 40.2 25.9 32.7 70.4 59.6 37.2 20.8 47.0 58.5 52.2 38.0 33.8 45.6 51.7 55.0 33.6 25.5 41.5
InternVL2.5 8 f 44.8 31.6 54.3 51.5 45.6 81.0 78.7 40.6 35.8 59.0 72.9 55.2 41.0 62.1 57.8 65.1 61.3 32.1 15.4 43.5
Qwen-VL2 8 f 41.7 40.4 42.7 32.8 39.1 79.3 69.2 43.4 20.8 53.2 61.9 55.2 32.0 51.4 50.1 50.7 56.1 31.4 23.7 40.5
Qwen-VL2.5 8 f 43.7 31.6 35.4 46.5 39.3 88.8 77.6 50.9 30.8 62.0 72.9 47.8 34.0 50.8 51.4 71.0 61.3 33.6 18.8 46.2
PLLaVA 16 f 25.0 24.3 36.6 23.2 26.5 33.5 54.3 29.9 23.3 35.3 30.5 32.8 34.0 14.7 28.0 50.7 37.4 29.9 21.8 35.0
ShareGPT4Video 16 f 36.5 30.9 40.9 10.1 29.6 33.0 40.4 24.5 43.3 35.3 45.8 40.3 39.0 16.9 35.5 44.1 31.0 34.3 25.0 33.6
Otter-V 16 f 15.6 25.7 26.2 21.2 21.9 30.2 22.3 20.8 22.5 24.0 22.9 17.9 20.0 23.1 21.0 38.2 26.5 33.6 27.6 31.5
VideoChat2-IT 16 f 25.0 30.9 39.6 25.3 30.2 20.1 47.9 11.2 26.7 26.5 42.4 47.8 34.0 24.9 37.3 43.4 27.7 31.4 23.0 31.4
LLaVA-Video 64 f 40.6 32.4 59.8 37.4 41.8 74.3 72.3 44.3 26.7 54.4 68.6 52.2 47.0 58.8 56.7 52.0 58.7 32.8 39.3 45.7
Video-LLaMA 8 f 25.0 25.9 11.5 19.7 18.9 28.4 33.2 23.9 20.0 26.4 39.5 32.5 23.2 15.4 27.7 40.1 26.6 33.1 26.2 31.5
VideoLLaMA2.1-AV 8 f 32.3 28.7 47.0 21.7 31.9 34.6 47.9 41.5 16.7 35.2 47.5 43.3 27.0 33.3 37.8 44.0 31.6 32.1 23.7 32.9
ImageBind-LLM 15 f 15.6 21.0 25.0 26.4 21.0 23.5 24.5 23.8 19.5 22.8 19.5 19.4 24.0 23.3 21.6 45.0 25.1 28.9 22.9 30.5
ChatBridge 4 f 27.1 15.6 30.2 14.4 21.8 31.4 41.5 23.7 8.7 26.3 31.4 23.9 41.0 16.9 28.3 43.7 25.3 30.4 24.4 30.9
OneLLM 15 f 26.0 29.4 34.1 21.1 26.1 29.0 38.3 20.6 27.8 28.9 32.2 38.8 23.0 22.4 29.1 43.4 26.4 29.5 23.2 30.6
GPT-4o 54.1 38.2 27.4 64.1 40.1 46.9 83.0 47.2 32.5 50.7 81.4 67.2 47.1 61.6 64.3 - - - - -
Gemini-1.5-Pro 57.3 54.4 53.0 67.2 55.9 87.1 77.7 52.8 71.7 72.3 78.0 65.7 54.0 75.1 68.2 90.1 76.8 66.4 84.6 79.5
Human 91.6 86.0 87.8 80.8 86.6 98.9 85.8 92.5 78.3 88.9 93.0 86.5 88.6 88.1 89.1 96.0 96.1 87.0 88.0 91.8
Table 1: The performance of SOTA video MLLMs on HumanVBench, grouped into 15 visual-only MLLMs, 5 audio-visual MLLMs and 2 proprietary ones. “Random” denotes random guessing, and “Human” indicates human-level performance. Task acronyms are defined in Figure 1. For each task, the best overall result is bolded, and the best open-source result is underlined. “-” means the model recognizes its lack of required capabilities for the task and thus refuses to answer. The results for all open-source models are averaged over five runs with different random seeds. Additional evaluation details are in the Appendix Section 9.
3.3.2 Question and Preliminary Answer Synthesis